OpenStudy (anonymous):

How do you find the measure of a central angle in degrees? The only information i have is circumference, area, diameter and radius of the circle.

OpenStudy (anonymous):

By area, do you mean the area of a sector of the circle? or the total area?

OpenStudy (anonymous):

the total area

OpenStudy (anonymous):

Hmm, that won't do. It doesn't look like you have any info you need to find the central angle.  You need to know something about the particular sector. The diameter is just twice the radius. The circumference and area of the circle only tell you about the whole circle, not the sector. By any chance, are you given a ratio of the circumference to the sector's arc length, or total area to sector area? That would be a step in the right direction.

OpenStudy (anonymous):

Well i am supposed to be finding a bunch of information on a ferris wheel (the singapore flyer). It asks me a bunch of questions in a row. Hold on ill post the questions.

OpenStudy (anonymous):

Complete the following tasks and calculations. You must show all work and steps to receive full credit. Name of the ferris wheel Diameter of the wheel (in meters) Number of cars or compartments Circumference of the wheel (in meters) Area of the wheel (in meters) Measure of a central angle in degrees Measure of a central angle in radians Arc length between two cars or compartments Area of a sector between two cars or compartments

OpenStudy (anonymous):

Okay, so there's some hidden info here. The compartments on a Ferris wheel have to be spaced evenly apart, right? How many compartments are there?

OpenStudy (anonymous):

28 compartments

OpenStudy (anonymous):

Are you there?

OpenStudy (anonymous):

28's a bit much to draw, so I'll just stick to 10 as an example:  Let's pretend the drawing is perfect, and that the circle can be cut into 10 equal-sized slices. This means that the central angle (the angle between adjacent cars) is \(\dfrac{360^\circ}{10}=36^\circ\), since there are 10 cars and thus 10 slices to the circle. Let \(r\) be the radius, then the circumference is \(2\pi r\). The arc length \(x\) of one of the 10 slices is given by the ratio \[\frac{2\pi r}{360^\circ}=\frac{x}{\frac{360^\circ}{10}}=\frac{x}{36^\circ}\]